A survey asks a random sample of 1500 adults in Ohio if they support an increase in the state sales tax from 5% to 6%, with the additional revenue going to education. Let p denote the proportion in the sample who say they support the increase. Suppose that 55% of all adults in Ohio support the increase. The st deviation of the sampling distribution is _______ 
Answer #1

7)

Solution

Given that,

p = 55% = 0.55

1 - p = 1 - 0.55 = 0.45

n = 1500

The standard deviation of the sampling distribution is ,  

=  p ( 1 - p ) / n

=   (0.55 * 0.45) / 1500 = 0.000165

= 0.0128

The standard deviation of the sampling distribution is 0.0128 .